OpenStudy (cicilybailey14):

I medal.... What is the common difference of the arithmetic sequence 9, -1, -11, -21, ...? -10 9 -9 10

OpenStudy (timmyk14):

-10 9 - 10 = -1 -1 - 10 = -11....